OpenStudy (anonymous):

HELP PLEASE ILL GIVE MEDALS Becca has a biscuit recipe that uses 2/3 of a tablespoon of salt for every 3/4 of a cup of flour. In order to have enough biscuits for her family that is visiting, Becca is making a batch of biscuits using 3 cups of flour and _______ tablespoons of salt.

OpenStudy (anonymous):

please help I'm not understanding this

OpenStudy (anonymous):

hello? plz someone HELP!

OpenStudy (fibonaccichick666):

ok so, this is a proportion you need to set up

OpenStudy (fibonaccichick666):

what you are saying is that the ratio will hold

OpenStudy (fibonaccichick666):

so we have to set up an eq, any ideas?

OpenStudy (anonymous):

umm...not really im terrible at math

OpenStudy (fibonaccichick666):

it's ok, and no one is terrible at math, they just haven't learned it yet. Don't get yourself down. so let me explain using different ex

OpenStudy (fibonaccichick666):

If for every 5 dollars I have I can buy 3 pairs of shoes, how many shoes can I buy with 15 dollars?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

9?

OpenStudy (fibonaccichick666):

yep 9

OpenStudy (fibonaccichick666):

so what you just solved was this proportion, \(\frac{5}{3}=\frac{15}{x}\)

OpenStudy (fibonaccichick666):

or dollars over quantity = dollars over quantity

OpenStudy (fibonaccichick666):

so, you solved for x, how did you do it

OpenStudy (anonymous):

i broke down 15 like this: 5, 10, 15. then i counted by 3's for each one and got 9(not a very good explanation)

OpenStudy (fibonaccichick666):

nope, that is another way of doing it, I want you to try it using this proportion though. It's a little bitt of algebra

OpenStudy (fibonaccichick666):

(a perfectly viable way too)

OpenStudy (fibonaccichick666):

do you know how to solve that eq? or would you like a walk through then another one you solve?

OpenStudy (anonymous):

explanation then another solve

OpenStudy (fibonaccichick666):

alright, so first, we are going to cross multiply to get rid of our fractions (and get our x on top. \[\frac{5}{3}=\frac{15}{x}\]\[5 x=3\times 15\]\[5x=45\] do you follow?

OpenStudy (anonymous):

yea

OpenStudy (fibonaccichick666):

so now we want to solve for x by dividing both sides by 5. \[5x=45\]\[x=\frac{45}{5}\] Then we simplify \[x=9\] do you see how I did that?

OpenStudy (anonymous):

yes that makes more sense thxs

OpenStudy (fibonaccichick666):

ok so now, how about you solve this one, \[\frac{4}{3}=\frac{x}{9}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x=12

OpenStudy (fibonaccichick666):

show me your steps before i tell you if you are right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i cross multiplied 4 and 9 and got 36 and i cross multiplied x and 3 to get 3x. then i divided 3 on both sides to get x by itself and then 36/3 is 12. so x=12

OpenStudy (fibonaccichick666):

very good, so now I want you to set up the proportion for your problem. do salt over flour=salt over flour

OpenStudy (anonymous):

ok but the question already went away srry i wasted your time.

OpenStudy (fibonaccichick666):

went away? you didn't waste it, you still can figure it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry i got to go. you helped a lot and i understand it thanks :)

OpenStudy (fibonaccichick666):

ohk? have a good one

OpenStudy (anonymous):

bye